### v1

# UNCOVERing the High-Redshift AGN Population Among Extreme UV Line Emitters

Abstract
JWST has revealed diverse new populations of high-redshift () AGN and extreme star-forming galaxies that challenge current models. In this paper, we use rest-frame UV emission-line diagnostics to identify AGN candidates and other exceptional ionizing sources, complementing previous studies predominantly focused on broad-line AGN. From a parent sample of 205 UNCOVER galaxies with NIRSpec/PRISM follow-up, we identify 12 C IV, He II, and C III] emitters. Leveraging the combined rest-optical and UV coverage of PRISM, we limit the emission-line model space using the sample’s [O III]/H distribution, significantly decreasing the overlap between AGN and star-formation models in the UV diagnostics. We then find that the five He II emitters are the strongest AGN candidates, with further support from two [Ne V] detections and one X-ray detection from Chandra. We cannot robustly quantify the AGN fraction in this sample, but we note that close to 20% of parent sample galaxies are AGN candidates. The lower-mass line emitters, which are consistent with both AGN and star-forming photoionization models, have more compact sizes and higher specific star formation rates than the parent sample. Higher-resolution and deeper data on these UV line emitters should provide much stronger constraints on the obscured AGN fraction at .
1 Introduction
Supermassive black holes (SMBHs) are ubiquitous in massive galaxies (Kormendy & Richstone, 1995; Richstone et al., 1998). Actively accreting SMBHs (i.e., active galactic nuclei, or AGN) provide an opportunity for detailed studies of accretion and shape our understanding of the co-evolution of SMBHs and their host galaxies (Ho, 2008; Haggard et al., 2010; Heckman & Best, 2014; Lacerda et al., 2020).
In the study of this co-evolution, high-redshift AGN are crucial, placing constraints on SMBH seeding (e.g., Volonteri & Reines, 2016; Dayal et al., 2019; Greene et al., 2020; Inayoshi et al., 2020; Bhowmick et al., 2022; Li et al., 2023; Zhang et al., 2023) as well as AGN feedback and their contribution to reionization (e.g., Fan et al., 2006; Madau & Haardt, 2015; Trebitsch et al., 2023; Dayal et al., 2024; Madau et al., 2024). However, the identification of high-redshift AGN is challenging, partially because low-metallicity stellar populations decrease the efficacy of traditional AGN line ratio diagnostics (e.g., Shapley et al., 2005; Groves et al., 2006; Kewley et al., 2013b, a; Juneau et al., 2014; Steidel et al., 2014; Feltre et al., 2016; Nakajima & Maiolino, 2022; Hirschmann et al., 2019, 2023).
Before the James Webb Space Telescope (JWST, Gardner et al., 2023), the most fruitful approaches to search for high-redshift AGN were UV, X-ray, and radio selection, each with their own biases (for a review, see Fan et al., 2023). From these searches, we knew of a couple hundred AGN at (e.g., Bañados et al., 2016; Wang et al., 2017). However, the handful of pre-JWST AGN all have broad lines (Mortlock et al., 2011; Bañados et al., 2018; Fan et al., 2019; Wang et al., 2021) and spectroscopic follow-up revealed rest-frame UV features comparable to those of the Vanden Berk et al. (2001) lower-redshift broad-line SDSS composite. We are therefore missing a census of the critical population of high-redshift narrow-line AGN. For these objects, which likely comprise the majority of the overall population, there is obscuration of the region of SMBH gravitational influence (i.e., the broad-line region), preventing observations of the AGN continuum and the gas velocity widths (e.g., Hickox & Alexander, 2018).
Selections of objects with strong rest-frame UV lines have yielded narrow-line AGN at cosmic noon (z23), where the Lyman break and rest-frame UV are observable with ground-based optical telescopes (e.g., Hainline, 2012; Alexandroff et al., 2013). These observations highlight the different UV features typical of the broad- and narrow-line regions, which extend to 1 pc and from 100 pc to 1 kpc, respectively (Hickox & Alexander, 2018). For successful selection of narrow-line AGN beyond cosmic noon, further work is necessary for the identification of effective diagnostics. Plaguing higher-redshift diagnostics are the uncertainties and degeneracies arising from the potential contribution of Wolf-Rayet stars, supernova-driven winds, shocks, and Population III stars (Hirschmann et al., 2019). As a result, studies of lower-redshift metal-poor galaxies are essential, as they serve as analogues for higher redshift star-forming contaminants with stellar ionization signatures that we could mistake for AGN (Erb et al., 2010; Stark et al., 2014; Mingozzi et al., 2023). These empirical approaches are complementary to theoretical efforts to separate AGN and star-forming photoionization models in parameter space, regardless of metallicity (e.g., Feltre et al., 2016; Gutkin et al., 2016; Hirschmann et al., 2019; Calabrò et al., 2023).
JWST is central to the efforts to identify more high-redshift AGN and constrain the trustworthiness of potential selection methods. Thanks to its wavelength range and unprecedented sensitivity, this facility has already both pushed the AGN redshift record above 10 (Bogdán et al., 2023; Goulding et al., 2023; Maiolino et al., 2024a) and provided a more representative sample of AGN. These AGN have been selected for their broad lines (Harikane et al., 2023; Kocevski et al., 2023a; Larson et al., 2023; Maiolino et al., 2023; Übler et al., 2023), color and morphology (Endsley et al., 2022; Onoue et al., 2023; Ono et al., 2023; Furtak et al., 2023a; Yang et al., 2023), and X-ray luminosity (Bogdán et al., 2023; Goulding et al., 2023; Kocevski et al., 2023b).
Unlike the previous record-holders, many JWST AGN are undetected in X-rays, UV-faint, and display significant reddening (e.g., Yue et al., 2024; Maiolino et al., 2024b). One common new population of AGN is the subset of so-called “Little Red Dots” (LRDs) confirmed to have broad lines (Furtak et al., 2024; Greene et al., 2024; Matthee et al., 2024; Wang et al., 2024a; Lin et al., 2024). LRDs are generally characterized by compactness and a continuum that is blue in the rest-frame UV and red in the rest-frame optical (Labbe et al., 2023; Barro et al., 2024; Akins et al., 2024; Kocevski et al., 2024; Kokorev et al., 2024).
Despite the successes of these selections, we know we are missing a critical population of narrow-line AGN. To this end, in conjunction with photoionization models, Scholtz et al. (2023) and Mazzolari et al. (2024) identified a total of 59 narrow-line AGN candidates using a combination of rest-frame UV and optical lines. In this paper, we apply the Feltre et al. (2016) UV diagnostics to UNCOVER (Bezanson et al., 2022) galaxies to better understand the narrow-line and UV-selected AGN populations at high-redshift. We highlight the power of the NIRSpec/PRISM data to identify and characterize line emitters, even when the low resolution blends many relevant lines. In particular, we leverage the rest-frame optical coverage to significantly cut down on the relevant star-forming models in the UV diagrams. We also use simulated PRISM spectra to demonstrate our ability to select high-EW AGN candidates.
The paper is organized as follows. In Section 2, we describe the JWST UNCOVER program, the PRISM reduction pipeline, and our parent sample of UNCOVER sources. We outline our line fitting procedures in Section 3. In Section 4, we present the UV diagnostic diagrams, their associated photoionization models, and the locations of literature star-forming galaxies, AGN, and ambiguous cases like GN-z11. In Section 5, we then place our sample of line emitters on these and other diagnostics to identify the subset of AGN candidates. The line emitters also have noteworthy host galaxy properties, as we explore in Section 6. In Section 7, we discuss the possible origins of the detected lines and conclude with an outlook on further possibilities for identifying and separating high-redshift narrow-line AGN from other sources of hard ionizing spectra.
2 UNCOVER Imaging and Spectroscopic Follow-Up
In this section, we provide an overview of UNCOVER (Bezanson et al., 2022) and the reduction of the PRISM spectroscopic follow-up used in this paper.
2.1 UNCOVER Overview
| MSA ID | ra [deg] | dec [deg] | F444W [mag] | log() | log(sSFR/) | [arcsec] | log([O III]/H) | ||
| 45209 | 3.555938 | 3.001 | 1.68 | 26.830.06 | |||||
| 36678 | 3.594799 | 3.004 | 1.83 | 25.510.02 | |||||
| 46973 | 3.572151 | 3.025 | 1.75 | 26.710.06 | … | ||||
| 11126 | 3.610193 | 3.047 | 2.09 | 28.680.18 | |||||
UNCOVER is a JWST Cycle 1 Treasury program consisting of ultradeep NIRCam imaging and spectroscopic follow-up of 600 sources (Price et al., 2024). The Cycle 1 imaging covers 45 arcmin2 in the CLASH/Hubble Frontier Fields z=0.308 cluster Abell 2744 (Lotz et al., 2017), where the boost from strong gravitational lensing allows for an effective depth of 31.5 rather than 30 mag. Much of the field also has medium band imaging from Cycle 2 (Suess et al., 2024). Both the lens model (Furtak et al., 2023b; Price et al., 2024) and the DR3 images and catalogs are publicly available (Weaver et al., 2023; Suess et al., 2024). Wang et al. (2024b) presents the prospector (Leja et al., 2017; Johnson et al., 2021; Wang et al., 2023) photometric redshifts and estimates of the physical properties.
Complementary to the JWST observations are multi-wavelength imaging campaigns of the UNCOVER field, including deep radio/sub-mm and X-ray observations that are critical to identifying AGN. ALMA 1.2 mm continuum photometry was extracted for all UNCOVER sources, with continuum RMS sensitivity 33Jy (DUALZ, Fujimoto et al., 2023a). The cluster also has deep (2.3 Ms) Chandra ACIS-I imaging (PI:A.Bogdán), which pushed the AGN redshift frontier with the discovery of a z = 10.1 AGN (Bogdán et al., 2023; Goulding et al., 2023).
Phase two of the UNCOVER program includes NIRSpec/PRISM follow-up of 600 sources of interest. With key science and legacy goals in mind, the extensive ranking of follow-up priorities included, for example, the highest-redshift candidates, dusty or quiescent galaxy candidates, and AGN candidates (Labbe et al., 2021). However, the majority of follow-up spectroscopic targets simply sample galaxies across a range of stellar mass and redshift. Price et al. (2024) detail the UNCOVER spectroscopic targeting strategy and spectroscopic data reduction.
2.2 NIRSpec/PRISM Data Reduction
Data reduction of the PRISM spectroscopy with msaexp (v0.6.10 Brammer, 2022) includes the following steps (Price et al., 2024). Starting with the MAST1 level2 products, msaexp corrects 1/f noise, masks snowballs and other artifacts, and subtracts bias from individual exposures. Next, the pipeline applies WCS, identifies the slits, flat-fields, and performs path-loss correction. On a common grid, 2D slits are extracted and drizzled. We then vertically shift and stack the 2D spectra before applying a local background subtraction. This extraction uses a Gaussian model with a free mean and width (Horne, 1986).
In order to account for the magnification of background galaxies due to strong gravitational lensing by the Abell 2744 galaxy cluster, we use the UNCOVER v2.0 strong lensing model (Furtak et al., 2023b), which was recently updated with 14 additional multiple image redshifts and 152 new cluster members (Price et al., 2024). The model was constructed with an updated version of the Zitrin et al. (2015) analytic code and includes five smooth cluster-scale dark matter halos and 552 cluster member galaxies. We refer the reader to Furtak et al. (2023b) for details on the model parametrization and optimization. The v2.0 model achieves an average lens plane image reproduction error of .
2.3 Parent Sample
The NIRSpec/PRISM spectra provide a wide range of emission lines () at low resolution (R50400). As such, for , the spectra include, at minimum, Ly through H. Because we require C IV for the Feltre et al. (2016) diagnostics (Section 5), we start with only the UNCOVER spectroscopic follow-up sources with (and redshift quality flag of 2 or 3 in Price et al. (2024)), leaving a parent sample of 205 galaxies. In Figure 1, we provide the distribution of magnitudes as a function of redshift for the parent sample and highlight the “Little Red Dots” (LRDs) from Greene et al. (2024) in addition to the line emitters that we present in Section 5. Table 1 includes key properties of the parent sample.
To obtain masses and star formation rates, we use the prospector (Leja et al., 2017; Johnson et al., 2021) fits to the NIRCam photometry, as outlined in Wang et al. (2024b). Because of the complex spectroscopic selection and the resulting bias of the parent sample, we also compare the masses and star formation rates to the 16,665 UNCOVER galaxies with . Wherever available, we incorporate the medium band photometry from the Cycle 2 program “Medium Bands, Mega Science” (Suess et al., 2024).
To measure the effective radius of each galaxy, we use pysersic (Pasha & Miller, 2023) to fit a single Sérsic profile to the F444W image, taken with a pixel scale of 0.04 arcseconds per pixel. We calculate uncertainties assuming the parameter posteriors are Gaussian, using the Laplace Approximation. After converting the radii in pixels to arcseconds, we use Planck Collaboration et al. (2020) cosmological parameters to obtain radii in kiloparsecs. We correct all these quantities with the magnifications.
3 Line Fitting
In this section, we describe our fitting approach for the UV lines used in the Feltre et al. (2016) diagrams (Section 3.2), H and [O III] (Section 3.3), H and [N II] (Section 3.4), [O II] and [Ne III] (Section 3.3), and the weak, high-ionization line [Ne V] (Section 3.5). We fit the UV lines, H and [O III], and H and [N II] of the entire parent sample but only fit the other lines for our final line emitter sample.
3.1 Modeling Choices
In each of the fitting regions, we use a power law continuum (normalized with the blue end of the fitting region) with a Gaussian added for each line. We require the lines within each fitting region to share an intrinsic width. We use flat priors on the power law index and amplitude between 3 and 3 and 1e22 and 1e18, respectively. For the fits to relatively weak lines (i.e., not for the +[O III] or +[N II] fits), we simultaneously fit for a multiplicative noise inflation term, using an allowed range of 0.1–10. A factor of 1.4 is typically a good fit. To find best fits and corresponding uncertainties, we use the PyMultiNest (Buchner et al., 2014) implementation of MultiNest (Feroz et al., 2009).
Before calculating likelihoods, we convolve the model with the instrument resolution, using the Jakobsen et al. (2022) JDOX pre-launch instrumental broadening. We adopt a conservative increased resolution of 1.3 (Greene et al., 2024). Point sources have better resolution by a factor closer to 2 (de Graaff et al., 2024; Slob et al., 2024), although this level is not preserved through the reduction and co-adding because the pixel size is large relative to the PSF. We note that lines have a systematic uncertainty on their FWHMs because of our use of this uniform inflation factor rather than a wavelength-dependent, size-dependent factor. To encapsulate this error, for the final emitters, we repeat the fits without any increase on the resolution. We propagate the resulting line flux differences into our plotted ratio and EW errors. This additional source of uncertainty is negligible for all but the highest-SNR sources (45092, 47584, and 45924). Before convolving the instrumental broadening and the model, we use a wavelength grid that oversamples the variable resolution by a factor of four. We then preserve flux while resampling onto the wavelength grid. In Section 5, we present the resulting diagnostic diagrams.
The line fitting presents several challenges: the spectral resolution is both low and variable (R50400), and our understanding of the line spread function (LSF) and noise is incomplete, including regarding the magnitude and impact of correlated noise. Once correlated noise has been quantified for UNCOVER, fitting should use a covariance matrix (as in JADES et al. in prep). To explore the robustness of deblending in light of these uncertainties, we test our fitting procedure on a series of mock spectra. We detail these tests in Appendix A but summarize our main conclusions here. Namely, we demonstrate that we can satisfactorily deblend lines and represent the corresponding flux uncertainties. However, we cannot rule out that a couple of our final sources could be false positives. For one source (34605), of mocks with no injected line yield 3 “detections” and we demonstrate that the false positive rates increase modestly when we use correlated noise in the generation of mocks (e.g., with a correlation coefficient of 0.6 between neighboring pixels).
In Figure 2, we show the spectra and corresponding fits for a few of the best sources in our final sample of rest-frame UV line emitters. We also emphasize the unmistakable presence of He II in several sources, regardless of the accuracy of the deblending procedure (Figure 3). In Figures 15–17, we show the rest-frame UV spectra for the remaining emitters, including marginal detections.
| MSA ID | log(C IV) | log(He II) | log(O III]) | log(N III]) | log(Si III]) | log(C III]) | C IV EW | He II EW | C III] EW | Fig. 7 Loc. | log([Ne III]) | log([O II]) |
| 34605 | AGN | |||||||||||
| 45092 | AGN | |||||||||||
| 47584 | AGN | |||||||||||
| 45924 | ||||||||||||
| 10646 2 | AGN | |||||||||||
| 42489 | … | |||||||||||
| 37005 | … | |||||||||||
| 17467 | … | |||||||||||
| 10646 1 | … | |||||||||||
| 10155 | … | |||||||||||
| 36755 | … | |||||||||||
| 23604 | … |
3.2 UV Lines
First, we fit C IV 1548, 1551, O III] 1661, 1666, N III] 1750, [Si III] 1883+Si III] 1892, and C III] 1907, 1909, with the fitting region extending to 0.02 rest-frame microns blueward of C IV and redward of C III]. For prior ranges on line fluxes, we use 1e21 up to double the LMFIT (Newville et al., 2016) maximum flux when fitting a power-law continuum plus C IV, He II, and C III]. We require all the lines to share a FWHM between 50 and 1000 km/s. Although the spectral resolution prevents us from constraining the line width, this range allows the fit to marginalize over this source of uncertainty. Because of the blending, we measure the EWs from the model fits.
Although we fix the redshifts for these initial fits to the overall solutions, we iterate on the fits for the final sample by freeing by 0.01 in each direction. This change allows a marginalization over wavelength calibration uncertainty. Three emitters (45924, 47584, 45092) have sufficiently high SNR to remove the noise inflation parameter in this updated fit. We therefore use the parameters from these free-redshift fits for the diagnostic plots and include the fluxes and EWs in Table 2.
3.3 H and [O III]
We also fit H and the [O III] doublet for the entire parent sample, with the doublet flux ratio fixed to 3.1 (Osterbrock, 1989). The fitting region includes 0.06 microns beyond the lines and we fix the redshift to the overall solution for the source. LMFIT once again provides an allowed range of values in the MultiNest fit. We only fit a shared narrow component to the lines. For the LRDs, which have broad H, we use the Greene et al. (2024) fits to account for both the narrow and broad components.
3.4 H and [N II]
To check for evidence of broad H, we fit H and the [N II] doublet, with a doublet ratio of 3.05 and the redshift allowed to vary by 0.01 in each direction. Following Greene et al. (2024), for the two-component fit to be preferred over the narrow-only one, we require a 3 improvement on and a 5 detection of the broad component by MultiNest. While there are no new broad-line AGN in our final emitter sample, we report one non-LRD BLAGN, which has evidence for UV lines at the 2.5 level (Figure 14).
3.5 [Ne V]
We fit the high-ionization line [Ne V] 3426 for the objects in the final line emitter sample. Because [Ne V] is situated in a complicated part of the continuum, we fit 0.04 microns (rest-frame) blueward of [Ne V] through 0.025 microns redward of the line, include He II 3204 in the model, and fix the redshift to the overall solution. Although the continuum near [Ne V] is not well-described by a power law, these choices allow for reasonable fits (Figure 2) and the two detections (45092 and 45924) are robust to precise fitting choices.
3.6 [O II] and [Ne III]
In our fits to [O II] 3726 and [Ne III] 3869, we use the continuum 0.04 microns blueward of [O II] and redward of [Ne III]. The model also includes H and H since they sit so close to [Ne III].
4 UV and Optical Diagnostic Diagrams
Now that we have the line strengths in hand, we provide an overview of the models that inform our use of UV emission-line ratios to classify galaxies. A common approach to identifying AGN is the use of ratios between nearby emission lines. These ratios probe the shape of the ionizing spectrum, which we expect to be harder in AGN than in star-forming galaxies (SFGs). For instance, in the local Universe, samples of AGN have been successfully selected at high values of [O III]/H and [N II]/H in so-called BPT diagrams (Baldwin et al., 1981; Veilleux & Osterbrock, 1987; Kauffmann et al., 2003; Kewley et al., 2006).
However, properties of the ISM also affect these line ratios. Modeling and observations both indicate that optical line ratios fail to separate AGN and SFGs beyond z1.5 because SFGs can move up to the AGN region and metal-poor AGN move towards the SF region of the diagrams (e.g., Shapley et al., 2005; Groves et al., 2006; Kewley et al., 2013b, a; Juneau et al., 2014; Steidel et al., 2014; Nakajima & Maiolino, 2022). This redshift evolution is driven by some combination of metallicity, abundance patterns, ionization parameter, sSFR, and electron density, all of which are significantly different at cosmic noon and higher redshifts (e.g., Brinchmann et al., 2008; Sanders et al., 2015, 2016; Hirschmann et al., 2017; Strom et al., 2017, 2018; Curti et al., 2024; Tacchella et al., 2024).
Nevertheless, the same models that predict the failure of optical line ratio diagrams reveal promising line ratios in the UV (e.g., Feltre et al., 2016; Hirschmann et al., 2023). It is also convenient to use UV ratios for high-redshift galaxies as rest-frame optical lines get pushed deeper into the IR. Moreover, some individual UV lines might be sufficient for AGN selection if the ionization potential requires an AGN.
In this section, we introduce the diagnostics we use to separate high-redshift AGN candidates from interesting star-forming galaxies. In Section 4.1, we provide an overview of these models as well as the simple cutoffs from Hirschmann et al. (2019). In Section 4.2, we leverage strong optical lines to limit the allowed model space and note the locations of literature AGN and SFGs with the necessary line detections. At the same time, these data underscore challenges on the modeling side, such as the inability to reproduce observed UV EWs in galaxies that do not seem to host AGN. Putting this together, we conclude this section with a prescription for AGN candidacy (Section 4.3). In Section 5, we then place our objects on ratio and EW diagnostics and note other relevant features of their spectra.
4.1 Models
There are a few UV diagnostics that we explore in this work, but, for the following reasons, we focus on the Feltre et al. (2016) UV diagnostic diagrams, which involve the ratios of He II 1640 with each of C IV 1548, 1551, O III] 1661, 1666, N III] 1750, [Si III] 1883+Si III] 1892, and C III] 1907,1909.
A proposed selection method that incorporates both UV and optical lines is the [O III]/H vs. [Ne III]/[O II] (OHNO) diagram (Zeimann et al., 2015; Backhaus et al., 2022; Larson et al., 2023; Cleri et al., 2023). However, the AGN region is contaminated with SFGs even at low-redshift (Backhaus et al., 2022) and sub-solar metallicities complicate both of the ratios in the OHNO diagram (e.g., Trouille et al., 2011; Scholtz et al., 2023). Nevertheless, some AGN may have ratios so extreme that they are indeed isolated in this space (e.g., Zeimann et al., 2015; Larson et al., 2023).
Individual lines can also serve as AGN indicators when their ionization potential is so high that stars are unlikely to account for any detection. The mere presence of [Ne V] in a galaxy spectrum is often considered sufficient evidence for an AGN, thanks to the ionization potential of 97 eV (e.g., Schmitt, 1998; Gilli et al., 2010; Mignoli et al., 2013), but there are cases of detected [Ne V] from seemingly inactive galaxies (e.g., Izotov et al., 2004, 2012, 2021). Especially at high redshift or low metallicity, [Ne V] may suffer from the same degeneracies as other UV diagnostics. However, the pairing of this line with [Ne III], for instance, helps to distinguish between AGN and other ionization sources (e.g., shocks or Pop III stars, Abel & Satyapal, 2008; Izotov et al., 2021; Cleri et al., 2023; Chisholm et al., 2024). Unfortunately, even in AGN, [Ne V] is intrinsically weak and highly reddened (e.g., Netzer, 1990), so we cannot rely on it for AGN selection, especially at low spectral resolution.
On the other hand, the Feltre et al. (2016) diagnostics are promising, even at the PRISM resolution. To identify these useful line ratios, Feltre et al. (2016) presented AGN models that use CLOUDY (Ferland, 1993; Ferland et al., 1998, 2013) and are parameterized by hydrogen density, ionization parameter, metallicity, dust-to-metal mass ratio of ionized gas, and the power-law index of ionizing spectrum. After pairing the AGN grids with the Gutkin et al. (2016) stellar models, Feltre et al. (2016) found that the key diagnostics leverage the expectation that relatively high values of He II 1640 signal contribution from an AGN narrow-line region (e.g., Villar-Martin et al., 1997; Allen et al., 1998). The success of He II as a tracer is linked to its recombination origin (with an ionization energy of 54.4 eV), whereas the nearby lines are largely collisionally-excited metal lines. We use all of the effective line ratios presented in Feltre et al. (2016) but emphasize the diagrams with C III]/He II vs. C IV/He II and C III]/He II vs. O III]/He II (e.g., Figure 6) because they include the lines that are typically easier to detect.
More recent studies have improved these models. As part of an effort to couple nebular emission lines to cosmological zoom-in hydrodynamical simulations, Hirschmann et al. (2019) included composite galaxies with both stellar and AGN contributions. By optimizing purity and completeness, they identified cutoffs between the AGN, composite, and SFG regions of the diagnostics, where the composite region includes galaxy models with a ratio of between black hole accretion rate and star formation rate. Since they incorporated updated spectra of Wolf-Rayet stars, we use their versions of the Gutkin et al. (2016) models in this work.
Still absent from these models, though, are rotation and binary products, which seem essential for understanding UV emission from star-forming galaxies (e.g., Eldridge & Stanway, 2012). To address this issue, Xiao et al. (2018) used CLOUDY but with BPASS111https://bpass.auckland.ac.nz/index.html models (Stanway & Eldridge, 2018). Crucially, they found that the resulting stellar model grids span the entire diagnostic space, potentially rendering the UV line ratios unhelpful after all. However, we demonstrate that the He II diagnostics remain empirically relevant for two reasons, as we discuss in Section 4.2.
Still, there are clear caveats to the models used herein. While the grids generally include wide parameter ranges, they mainly stick to solar abundance ratios and the Feltre et al. (2016) AGN models only scale the ionizing spectrum with accretion luminosity. The currently inevitable uncertainty, though, is that there are few observations with which to check and calibrate the modeled UV emission (e.g., Whitler et al., 2023). For instance, metal-poor models (i.e., ) have not been sufficiently tested with observations since there are limited studies of resolved lower-Z O stars (e.g., Telford et al., 2023).
4.2 Semi-Empirical Diagnostic Boundaries
4.2.1 [O III]/H Limits Model Grid Overlap
There is significant overlap between the full AGN and SFG model grids, but we can directly cut down on the grid parameters using other information offered by the PRISM wavelength coverage. In particular, after cutting out models with [O III]/H values inconsistent with the distribution within our final sample, we see renewed support for the usefulness of the diagnostics.
At these redshifts, we no longer trust AGN selection methods hinging on optical ratios like [O III]/H, because of the sensitivity to not just the ionizing spectrum but to the ionization parameter and metallicity. However, the sensitivity to these parameters allows [O III]/H to constrain them and thus to decrease the degeneracy between possible origins of the observed UV line ratios.
We plot the [O III]/H distributions for our parent sample, our line emitters, and the SFG model grids in Figure 4. Using the minimum [O III]/H of 3.0 found among our emitters, we apply a cut to the model grids. In Figure 5 we show the model diagnostic diagram before and after applying this cut. This cut drastically reduces the span of the restricted SFG model grid, such that there is limited overlap between the AGN and SFG models.
For each metallicity and ionization parameter combination, we also plot the fraction of models still included after the cut. The extreme metallicities in the model grids are then disfavored or entirely removed, with the highest fraction of models kept at Z0.005. The effect on ionization parameter, on the other hand, is monotonic: for the stellar models, values of logU lower than cannot survive this cut, and the cut’s effect decreases when moving to higher ionization parameters in the grid. Interestingly, although the full Xiao et al. (2018) model grid extends more than the Gutkin et al. (2016) one into the composite and AGN regions, it occupies less parameter space than the Gutkin et al. (2016) models once the cut has been applied to both. We adopt the updated Gutkin et al. (2016) models in the plotted diagnostics to show the more conservative span of SFG models.
4.2.2 Observed Line Ratio Diagram Samples from the Literature
Literature objects also support the viability of the UV diagnostics. In Figure 6, we place literature AGN and SFGs on the two most important ratio diagnostics (C III]/He II vs. C IV/He II and C III]/He II vs. O III]/He II). Here we explain the basics of the included objects and highlight key take-aways that affect how we interpret the locations of the UNCOVER galaxies. In short, most literature AGN and SFGs with the necessary line detections live in the Hirschmann et al. (2019) composite region, but narrow-line AGN and SFGs occupy different areas within that region.
First, we note that these diagnostics only apply to the narrow-line region of AGN. The Feltre et al. (2016) AGN models do not consider the temperature and density of the broad-line region and we see that the broad lines of SDSS quasars are deep in the star-forming region. We show this discrepancy using both the SDSS composite (Vanden Berk et al., 2001) and the subset of nitrogen emitters (Batra & Baldwin, 2014). At PRISM resolution in the rest-frame UV, we generally cannot distinguish broad from narrow lines, so we do not expect to select any AGN with UV features dominated by the broad-line region.
Second, the diagnostic works for a variety of narrow-line AGN populations (UV-selected at cosmic noon (Hainline et al., 2011; Alexandroff et al., 2013), optically-selected locally (Nagao et al., 2006), X-ray-selected (Nagao et al., 2006), selected within a sample of C III] emitters (Le Fèvre et al., 2019), and radio galaxies (De Breuck et al., 2000)), but AGN often live in an ambiguous region with model overlap or even just at the edge of the SFG models. Similarly, the star-forming galaxies tend to be consistent with models, but with most being in the Hirschmann et al. (2019) composite region. We include star-forming galaxies and stacks as well as their local low-metallicity dwarf analogues. Despite the benefit of the low-redshift analogues for probing poorly-understood low-metallicity star formation, the few seemingly star-forming objects sitting among the AGN in Figure 6 are all higher-redshift. However, for the star-forming galaxy most consistent with the literature AGN in the log(C III]/He II) vs. log(O III]/He II) diagnostic, the C IV detection is marginal and it is difficult to rule out contribution from an AGN (Amorín et al., 2017). Because of these ambiguities and the PRISM spectral resolution, we focus on identifying the best AGN candidates, noting that these objects would benefit from higher-resolution follow-up.
4.3 AGN Candidacy
We have seen that the (post-cut) models and literature sources support the utility of the UV diagnostic diagrams for the identification of narrow-line AGN candidates. Still, there is significant overlap in the diagrams thanks to different possible physical causes of the ionization in the rest-frame UV.
To help break the remaining degeneracy, the existing data suggest we can leverage C IV information and EWs. Stellar systems with relatively strong He II, which could be misinterpreted as AGN in the UV diagnostics, often have a broad He II component and a P Cygni or weak C IV profile (Conti et al., 1996; Leitherer et al., 1996; Shapley et al., 2003; Erb et al., 2010; Sobral et al., 2015; Nanayakkara et al., 2019; Saxena et al., 2020). This combination of features is usually attributed to Wolf-Rayet stars, unless the EWs are above 3, in which case Very Massive Stars () are invoked (e.g., Gräfener & Vink, 2015; Martins et al., 2023; Wofford et al., 2023; Upadhyaya et al., 2024). Even with this absence of velocity information, we can be more confident in AGN candidates if they have C IV emission.
EW measurements present another potential way to discriminate between different ionization mechanisms. In the C III] EW vs. C III/He II and C IV EW vs. C IV/He II planes, Nakajima et al. (2018) suggested cutoffs in which high EW alone can require an AGN. For example, their modeling indicates that objects with C III] EW above 20 must be AGN. Complicating this approach is the fact that star formation models fail to account for the highest observed EWs of C IV, He II, and the nearby metal lines (e.g., Shirazi & Brinchmann, 2012; Jaskot & Ravindranath, 2016; Steidel et al., 2016; Senchyna et al., 2017; Berg et al., 2018; Nanayakkara et al., 2019; Plat et al., 2019; Olivier et al., 2022; Saxena et al., 2020, 2022; Stanway & Eldridge, 2019). Particularly because this problem is clearest at low metallicity (e.g., Plat et al., 2019), there is no consensus as to whether AGN alone bridge this gap (e.g., Shirazi & Brinchmann, 2012). The only AGN-free models that have replicated observations require bursty star formation histories, X-ray binaries, radiative shocks, and a top-heavy IMF (Lecroq et al., 2024).
Thus, degeneracy in these diagnostics will persist at least until we have better emission-line models for a variety of ISM conditions. At our spectral resolution, we can therefore require our most confident AGN candidates to be consistent with the AGN models in the line ratio diagnostics and to have C IV detected in emission. We can further boost confidence in an AGN classification for cases with relatively high EWs, but it remains to be seen the extent to which such EWs can be modeled by other processes. Still, especially as photoionization modeling evolves to describe larger data sets, we cannot rule out that other objects may also host AGN.
5 Diagnostics with UNCOVER UV Line Emitters
In this section, we introduce the sample of UV line emitters and identify the subset that we consider AGN candidates based on their location in the UV diagnostic diagrams. To do so, we place these objects on He II-based ratio (Section 5.1) and EW-ratio diagnostics (Section 5.2). We discuss evidence for broad H in one non-LRD source (Section 5.3) and then highlight the potential presence of [Ne V] in a couple of sources (Section 5.4). We also place this sample on the OHNO diagram (Section 5.5). These efforts divide the sample of line emitters into AGN candidates and ambiguous sources that may be AGN or exceptional star-formers (Section 5.6).
5.1 UV Ratio-Ratio Diagnostics
Our final sample of UV line emitters consists of the galaxies with a detection above 5 for C IV or above 3 for He II or C III]. We choose these lines since He II is in every ratio in the diagnostics, C III] is in every y-axis ratio, and C IV (non-)detections are of particular note as we try to separate sources of ionization (Section 4.2.2). Since C IV is unblended, it is easier for a noise feature to be deemed significant at the 3 level; a comparable feature at the location of He II or C III] would split flux with O III] or Si III] in the fit. We therefore use the higher threshold for C IV detections, finding that our suite of mocks also supports this choice (Section A). In Figures 15–17, we show the fixed-redshift and free-redshift fits to this final sample of line emitters. We use the free-redshift fits for the diagnostics to marginalize over uncertainty in the redshift and wavelength solution. From the parent sample of 205 galaxies, we have five He II emitters, eight C III] emitters (four of which are also He II emitters), and seven C IV emitters. Of the C IV emitters, three have only C IV detected.
In Figure 7, we place these line emitters on four diagnostics. Each diagnostic involves three lines, requiring most sources to be plotted with 3 limits. In cases with only C IV or C III] detected, we fix the limits at a y- or x-axis value of one, respectively. These C IV and C III] emitters without He II detected are consistent with both the AGN and SFG model grids. We therefore focus our discussion of AGN candidates on the He II emitters, which, apart from 45924, are more consistent with the AGN models in the C III]/He II vs. C IV/He II diagram and in an ambiguous region in the remaining diagrams. Because of the locations of literature AGN in these diagnostics (Figure 4.2.2), we do not require our candidates to sit in the Hirschmann et al. (2019) AGN region.
Using the MSA IDs, we label the AGN candidates. LRD 45924 is deep in the SFG model region, as is expected for broad-line AGN (Section 4.2.2). One candidate (47584) has particularly low C IV/He II, making it inconsistent with both model grids. A complicated underlying C IV profile (i.e., with P-Cygni, nebular, and resonant contributions) may drive this discrepancy with the models, since the PRISM spectral resolution requires us to fit all of these components with a single Gaussian.
The list of narrow-line AGN candidates includes one of the two seemingly resolved (Weaver et al. in prep) galaxies in an ionization bubble at (10646, Fujimoto et al., 2023b), a source at a very similar redshift to LRD 45924 (47584), the most massive galaxy in the parent sample (45092, Bezanson et al. in prep), and the one He II emitter without C III] also detected (34605). In Appendix B, we comment on interesting aspects of each of these sources and of the broad-line candidates.
5.2 UV EW-Ratio Diagnostics
Similar EW-based diagnostics are helpful since they only examine two emission lines at a time and allow us to compare the EWs to those of literature AGN and star-formers.
We include the C IV- and C III]-based EW diagnostics in Figure 8. We mark the Hirschmann et al. (2019) cutoffs between AGN, composites, and SFGs, as well as the Nakajima et al. (2018) version, which is just between AGN and SFGs. The Nakajima et al. (2018) separation requires anything at relatively high EW to have an AGN contribution, and thus places in the AGN region supposed high-redshift, extreme star-formers observed with JWST (Topping et al., 2024a; Castellano et al., 2024). Extreme UV EWs are not restricted to a couple outlier high-redshift galaxies; even composite PRISM spectra show a correlation between C III] EW and redshift (Roberts-Borsani et al., 2024). There is thus better agreement between the Hirschmann et al. (2019) cutoff and the included literature populations.
Turning to the locations of the objects in our sample within the diagnostic plots, we see that besides 45924, the He II emitters are consistent with the AGN region, providing important additional support to the findings from Figure 7. Still, ambiguity remains, at least for the objects with both lines detected, thanks to the proximity to some of the literature galaxies deemed SFGs.
Furthermore, we present a few C IV and CIII] emitters (45092, 45924, 36755, 37005, and 23604) with comparable EWs to those observed in exceptional objects GN-z11 (Bunker et al., 2023), RXCJ2248-ID (Topping et al., 2024a), and GLASS-z12 (Castellano et al., 2024). In fact, every He II emitter in our sample has a He II EW rivaling the one measured for GN-z11 in Bunker et al. (2023). Thus, these diagnostics lend support to the AGN candidacy and demonstrate the remarkable nature of the UV EWs seen in this sample.
5.3 Broad H
We find evidence of broad H (FWHM1000km/s) for one non-LRD. Roberts-Borsani et al. (2024) also recognized this object (11254, z=6.87) as a broad-line AGN candidate. We show the narrow-only and two component fits and their resulting residuals in Figure 14. We also include the UV region because of the evidence for a few of the fitted lines.
5.4 [Ne V]
We detect [Ne V] above 3 for LRD 45924 and for narrow-line AGN candidate 45092. None of the 3 limits for the other emitters come within a factor of 10 of the fitted flux for 45092, whereas three upper limits are consistent with the measured flux for 45924. Both objects are then consistent with the AGN and composite regions (but not the SFG or Pop III/IMBH ones) of the [O III]/ vs. [Ne V]/[Ne III] diagnostic (Cleri et al., 2023).
5.5 OHNO
In Figure 9 we compare the relative locations of our objects and several literature sources on the OHNO diagram. Consistent with other high-redshift and low-metallicity studies, we see that all of our objects are above the Backhaus et al. (2022) cutoff, implying they are selected as AGN. We confirm that this diagram does not reliably separate AGN and SFGs. Lastly, while we plot very few literature SFGs, we do see that the star-formation–powered sources do not go to lower [O III]/H than the He II and C III] emitters, so the model grid cut also applies to them.
5.6 Characterization of Our Line Emitters
These fits and diagnostics leave us with four narrow-line AGN candidates, two broad-line AGN, and seven objects requiring higher-resolution data for classification. In particular, we designate the five He II emitters (Figure 15) as AGN candidates. Only one of them sits on the SFG model grids, but that is to be expected considering it is a known LRD with Balmer line widths of 4500 km/s (Greene et al., 2024). To this group we can add 11254, an object with a broad H component and evidence of C IV, O III], and C III] emission.
The remaining seven emitters (Figures 16 and 17) lack He II detections, precluding strong AGN candidacy with this approach. However, they are all consistent with both the AGN and SFG models in the UV diagnostics. These C III] and C IV emitters are either AGN or exceptional star-formers with EWs that further challenge models of SFGs.
We provide individual descriptions of the AGN candidates in Appendix B, but note here that of the narrow-line AGN candidates, one (45092) has clear detections of both [Ne V] and X-ray emission, two (45092 and component 2 of 10646) present C IV emission, and two (45092 and 47584) have evidence of N III]. Like 45924, the strongest narrow-line AGN candidate (45092) also has N IV]. We must consider that candidates without C IV detections, which we consider to be lower-confidence anyway (Section 4.3), may be compact, GN-z11-like objects, with metal-poor (but nitrogen-enhanced) bursty star-formation (Bunker et al., 2023; Marques-Chaves et al., 2023).
6 Galaxy Properties
In this section, we comment on the host galaxy properties of the line emitters, both candidate AGN and exceptional star-formers. The opportunity to derive these properties is another benefit of the extensive wavelength coverage at our disposal with JWST. In the presence of a narrow-line AGN, it is reasonable to assume that the continuum is galaxy-dominated (e.g., Hickox & Alexander, 2018), but higher-resolution data could determine the AGN contribution to the photometry. We present these host galaxy properties in three figures and, in each case, compare the emitter sample(s) to the parent sample of UNCOVER galaxies with spectroscopic follow-up. We note that the parent sample itself involves a complex selection function (Price et al., 2024). As such, for specific star formation rate (sSFR) and stellar mass, we also call attention to the UNCOVER photometric sample of galaxies with . Wherever available, these photometric redshifts and properties incorporate the full medium band coverage provided by Cycle 2 program “Medium Bands, Mega Science” (Suess et al., 2024). In Table 3, we list the , , and percentiles of these properties for the parent sample, emitter sample, and emitter subsamples.
Here, we explain the three figures and overall differences between the parent sample and the sample of 12 emitters. In Sections 6.1–6.3, we then highlight commonalities within each subsample of line emitters. First, we mark the final sample of emitters in magnitude vs. redshift space and include the Greene et al. (2024) LRDs to compare to the locations of known broad-line AGN (Figure 1). The line emitters span the parent sample range in both axes but are systematically brighter.
In Figure 10, we plot sSFR averaged over 100 Myr vs. stellar mass for the sample of line emitters, with contours for the photometric and spectroscopic samples. Almost all the emitters are at high mass and/or high sSFR relative to the parent sample, with none within the one contour of the sample. In particular, the AGN candidates are clearly biased to high stellar masses, while the other emitters are lower mass than the AGN candidates but present modestly higher sSFRs than typical in the parent sample.
Lastly, we compare stellar masses to effective radii in Figure 11 and compare to the spectroscopic parent sample. A handful of emitters are remarkably compact, especially for their mass, resulting in statistical significance to the size difference for the emitters and the parent sample. We now discuss the locations of the subsamples of emitters with He II, C III], and C IV detected.
6.1 AGN Candidates (He II Emitters)
The He II emitters are the clearest AGN candidates in the ratio-ratio and EW-ratio diagnostics (Figures 7 and 8), or, in the case of 45924, known broad-line AGN. These objects are all relatively bright and massive, with these differences being the most significant among the comparisons in this section.
The He II emitters are all brighter than mag in F444W. Relatedly, they are more massive than the other emitters and more massive than is typical in the parent spectroscopic and photometric samples. The strongest narrow-line AGN candidate (45092) has the highest mass in the parent sample and sits at low sSFR, as it is a post-starburst galaxy (Bezanson et al. in prep). The closest parent sample object to 45092 in sSFR-mass space is 18407, a massive () quiescent galaxy at (Setton et al., 2024). On the other hand, a couple of the other exceptionally high-mass He II emitters are at high sSFR.
Because this group is so massive, there are few comparable galaxies for size comparison. Still, they seem compact for their masses, although that observation is not statistically significant. The least compact AGN candidates are 34605 and 47584, which also lack C IV detections. On the other hand, for either F444W magnitude or mass, a two-sample Kolmogorov-Smirnov (KS) test comparing the He II emitters with the parent sample shows statistical significance (p-value 0.05). This result implies that the He II mass and magnitude distributions are different from those of the parent sample; this conclusion is also clear in Table 3. We further address this mass bias in Section 7.
6.2 C III] Emitters
Of the subsamples, the C III] emitters have the closest distributions to the parent sample, particularly if we exclude the four that are also He II emitters.
The C III] emitters are also under-represented in number at the faintest magnitudes, but their magnitudes do range from 21–28. As is apparent in Figure 10 or a KS test, there is not sufficient evidence to conclude that these emitters have different sSFR or mass distributions than the parent sample. However, if we exclude the He II emitters from this calculation, the remaining C III] emitters do have significantly higher sSFR. C III] emitter 17467 is the galaxy with the highest sSFR in the sample. We also note that it is magnified by a factor of 133, or 25 times more than the next most magnified source in the emitter sample.
C III] emitters are, on average, more compact compared the parent sample, although a KS test does not allow us to reject that their sizes are drawn from the same distribution. Still, broad-line AGN 45924 is more compact than the other line emitters, and this mean difference holds up whether or not we include it in the average for this subsample.
6.3 C IV Emitters
The C IV emitters are the most compact subsample; the five most compact emitters are all C IV emitters. They are also potentially over-represented at the highest redshifts: three of the seven C IV emitters are above redshift 6, while this is only true for 22% of the parent sample. Furthermore, in the absence of the AGN candidates, they are typical in mass and high in sSFR. Although the C IV emitters do not generally skew to higher masses, the two most massive emitters are precisely the He II emitters with C IV also detected.
| Parent (205) | Emitters (12) | He II (5) | C III] (8) | C IV (7) | |
| F444W mag | |||||
| z | |||||
| log(100 Myr sSFR / ) | |||||
| [kpc] | |||||
| log([O III]/) |
7 Discussion and Summary
We have demonstrated the usefulness of NIRSpec/PRISM for the identification of high-redshift (e.g., ) narrow-line AGN candidates. Central to these efforts are careful deblending and the use of the sample’s [O III]/H to cut down on the relevant SFG models. The role of this cut underscores the essential constraining power of the PRISM wavelength range; no JWST medium-resolution gratings could capture both sets of lines in one integration.
We select 12 strong C IV1548,1551, He II1640, and CIII]1907,1909 emitters from a sample of 205 galaxies. One He II emitter (45924) is a known LRD while the other four are the most promising narrow-line AGN candidates. We identify one additional broad-line AGN (11254, Figure 14). If the remaining seven emitters are indeed SFGs, their UV EWs largely require updated models.
Here, we discuss the implications of the line emitter properties, contextualize the AGN candidates using other JWST studies, and highlight opportunities to expand on this work through higher-resolution data.
7.1 NLAGN Candidates
The He II emitters, which include the best narrow-line AGN candidates, are brighter and higher-mass than the other emitters and are outliers in the parent sample; 17% of galaxies in the parent sample with are also AGN candidates. If we assume we are unbiased for the highest-mass hosts, it can be compared to the Maiolino et al. (2023) finding that 10% of galaxies host broad-line AGN with .
The other galaxies in the line emitter sample have 3 line ratio limits consistent with both model grids. As a result, the AGN candidate separation could be a selection effect in which we are insufficiently sensitive to detect He II in lower-mass objects, because they likely host lower-mass SMBHs, which would have weaker emission lines if they are otherwise comparable.
Comparing to other efforts focused on narrow-line AGN candidates, our findings are not inconsistent with rates found in JADES (Scholtz et al., 2023), although we are hesitant to quantify an exact AGN fraction because of our spectroscopic selection function. In particular, Scholtz et al. (2023) presented 27 NLAGN candidates at . Most candidates required a conservative version of optical diagnostic diagrams for selection and higher-resolution spectra yielded the six candidates identified with C IV, C III], He II. The clearest difference is the candidate mass range: none of their candidates have , further supporting the possibility that we can only detect the highest-mass AGN with PRISM and rest-frame UV diagnostics.
7.2 C IV emitters
The C IV emitters seem to include AGN and likely star-formers, but trace a potentially different population from the C III] emitters. Three C IV emitters are He II emitters and AGN candidates, while the other four are consistent with either population but all at high sSFR. Key properties potentially setting C IV emitters apart are their compactness, their relatively high redshift distribution, and their lower [O III]/H. However, with the exception of compactness, this conclusion is not robust with the current sample size (Table 3). Additionally, 36755 and 23604 boast comparable C IV EWs ( and , respectively) to the exceptional galaxies investigated in Topping et al. (2024a) and Castellano et al. (2024).
C IV emission is seen in a number of contexts: AGN (Figure 6), low-metallicity sources (Stark et al., 2015; Mainali et al., 2017; Schmidt et al., 2017; Witstok et al., 2021), and Lyman continuum emitting galaxies (e.g., Schaerer et al., 2022). It is unclear the extent to which properties like metallicity, SFR, ionization parameter, and ISM density play a role in the presence and strength of C IV emission. In the case of C IV-emitting star-forming galaxies, metal-poor gas and young stellar populations seem essential (Senchyna et al., 2017; Topping et al., 2024b); these requirements may explain the high redshift and relatively low [O III]/H of our C IV emitters. However, the JADES Ly emitter stack has higher C IV EW than the higher-redshift one (Kumari et al., 2024), and real redshift evolution is difficult to establish definitively given the changing spectral resolution.
Lastly, we note an apparent connection to Ly that is consistent with literature findings. C IV emission can be a tracer for the escape of Lyman continuum photons (e.g., Naidu et al., 2022; Mascia et al., 2023) and among faint Ly emitters, Feltre et al. (2020) only detected C IV in emission in their stack with the highest Ly EWs (i.e., ). This discrepancy is likewise apparent in the median spectra of LAEs and non-LAEs across public NIRSpec PRISM data, despite the stacks having the same average redshift (Roberts-Borsani et al., 2024). In our sample of line emitters, high EW Ly emission is common among the C IV emitters but not in the other subsamples. However, detailed quantification of this difference is beyond the scope of this work.
7.3 The Future of NLAGN with JWST
We have demonstrated both the opportunities and obstacles for identifying NLAGN using JWST. Updated models are necessary because of the unprecedented EWs presented here, but there is also more to do with existing models. Although we illustrated the usefulness of the wavelength range, we have not compared AGN vs. SFG model fits to the full PRISM spectra. Additional constraints will come from the continuum shapes and numerous other detectable emission lines unexplored in this work. By modeling the full observed wavelength range, one can also isolate which features we cannot explain with self-consistent modeling (e.g., Cameron et al., 2024; Tacchella et al., 2024; Vidal-García et al., 2024). Increasingly clear is the necessity of including stellar models with non-solar abundance ratios (e.g., Grasha et al., 2021) and new binary models that can reproduce the observed He II emission in known SFGs (e.g., Lecroq et al., 2024). Another promising approach is the use of a flexible, source-agnostic tool that can accurately reconstruct the ionizing spectrum and whether binary stellar models are necessary for its production (Cue, Li et al., 2024).
Larger samples and higher spectral resolution would make the use of higher-ionization lines more feasible. Although this approach is necessary, the role of PRISM in these efforts will persist, because of its ability to efficiently observe a broad wavelength range. For instance, a 7 [Ne V] detection in a z=5.59 galaxy required NIRSpec high-resolution and an integration time of 14.7 hours (Chisholm et al., 2024). The [Ne V] flux was 24 times less than that of H, with a rest-frame EW of 11. The EW and strength relative to other emission lines are significantly higher than those found in star-forming galaxies with [Ne V] detections (e.g., Izotov et al., 2004, 2012, 2021), supporting the case for the line’s AGN origin. Only four of our objects have 3 upper limits ruling out that they have at least the same [Ne V] flux as this object. Only one object in the Scholtz et al. (2023) JADES sample boasted a [Ne V] detection and He II and C III] were not detected in this object, making it difficult to pinpoint relative to AGN and SFG models.
We encourage additional multi-wavelength and high-resolution follow-up of the AGN candidates and exceptional star-formers presented here; among the potential SFGs, we emphasize MSA IDs 37005, 36755, and 23604 because of their unprecedented UV EWs and 17467 because of its low mass, high sSFR, and the extent of its magnification. There is also abundant PRISM data that is unexplored with this approach (Roberts-Borsani et al., 2024). Although the PRISM rest-frame UV selection method presented here comes with myriad challenges, it has the potential to uncover poorly-understood high-redshift NLAGN and metal-poor populations, especially with the help of modeling advancements and higher-resolution follow-up.
Appendix A Fitting Simulated Rest-Frame UV Spectra
To evaluate the efficacy and limits of the MultiNest fitting of the UV lines (Section 3), we analyze mock versions of our final line emitters. Here, we describe the set-up for these simulated spectra and the key take-aways that inform our analysis of the MultiNest fits to the emitters. In short, MultiNest can recover injected flux values and provide reliable uncertainties. Furthermore, we confirm that the presence of correlated noise affects these mock results, increasing the chance of false detections at low SNR. However, these modest contributions do not significantly change our interpretation of the results. Although our fitting approach can deblend lines sufficiently well for our purposes, blended line ratios can be more biased than individual fluxes, since a slight underestimation of one line frequently corresponds to an overestimation of the other component of the blend. However, such a bias may be restricted to ratios that include a non-detection. We also highlight the few line emitters that could be false positives, albeit with low probability.
As the AGN candidates, the He II emitters constitute the most important group in this work. They also present the most difficult deblending problem since He II is included in all the diagnostics. C III] is blended with Si III], but C III] should be stronger than Si III] and the strength of either of those lines relative to He II bolsters the case for star formation as the origin of the ionization. We therefore perform MultiNest fits to many mocks of each He II emitter to check our confidence in the He II detections. This approach addresses a few potential concerns, with the key one being a check of the likelihood that we are attributing significant He II flux to what is really O III].
First, for each He II emitter, we use MultiNest to fit 100 mock spectra with parameters equal to the best fit ones for the real spectrum. For each iteration, we add Gaussian noise to each spectral bin. For the standard deviation of this noise, we use the true error array scaled by the best-fit multiplicative error scaling. In the case of 45092, 45924, and 47584, we use the original error array and mirror the original fits by not including an error scale factor in the model of the mock spectra.
We then repeat this test but with the injected He II equal to the 3 lower limit from the original fit to the data and the O III] adjusted such that the total blend flux remains the same. Through this set-up, we confirm that these kind of underlying flux values do not yield the best-fit He II values that we found with the real data. In other words, we do not need to be concerned that our He II detections are actually O III] or noise. Figure 12 summarizes the findings from these two tests. Namely, for our purposes, MultiNest successfully identifies the He II flux values and the associated uncertainties. Moreover, none of the fits to the He II 3 lower limits overlap with the best fit He II values, which we interpret as indication that we have not falsely attributed He II flux to O III] or noise.
A slight modification on this approach also allows us to understand the effect of correlated noise, which is surely present but has not yet been robustly quantified in UNCOVER. Once available, one should work with a covariance matrix when fitting (as in JADES et al. in prep). In particular, we create mocks with the injected He II at the 3 lower limit but construct the added noise to have correlations between neighboring pixels. We perform this test with correlation coefficients of 0.3 and 0.6, finding that only low SNR cases (i.e., 34605) present a clear trend with increasing correlation coefficient. Namely, while 1% of uncorrelated noise mocks yield fitted fluxes above the real spectrum’s best-fit He II value, this rate increases to 2% and 4% for correlation coefficients of 0.3 and 0.6, respectively. Regardless of the specifics of the noise, it is clear that He II emitter 34605 has some chance of being a false positive.
We also examine potential bias in the ratio of O III]/He II, which is one of the key diagnostic ratios. Here, we have more cause for concern, since the uncertainties in each line are correlated, leading to systematic bias in the ratio. First, our 3 O III]/He II upper limits for sources without O III] detections may be unnecessarily high, since the fits attribute too much flux to the undetected O III] line. In particular, as shown in Figure 13, for the source with the lowest injected O III] flux and O III]/He II (45092), the mock fits are biased towards higher O III]/He II ratios. The median ratios from the mocks are just outside the 1 ratio uncertainties from the best fit to the data, but we indicate these ratios as upper limits in the diagnostics, anyway. This bias simply indicates that the true O III]/He II ratios are likely lower than the upper limits indicate, making the sources even clearer AGN candidates.
On the other hand, the O III]/HeII ratios of 47584 are under-estimated. Thus, the O III] flux may drive the direction of the bias, since 47584 had the highest injected O III] flux, as well as the most over-estimated He II flux (Figure 12). This bias is slight, though, and such an offset would be inconsequential on the diagnostics. However, for the one He II emitter with O III] also detected (45924), the MultiNest and mock O III]/He II ratios are in good agreement.
These mocks also allow us to test the redshift fitting in this region. The redshift fits work well, with a mean deviation from the injected value of 0.002. There are rarely any fits close to the 0.01 prior bounds. Furthermore, the success of the redshift fitting does not rely on a strong C IV flux, despite C IV being one of the two unblended lines and the only one to sometimes be strong in the real spectra.
Lastly, for all 12 line emitters, we perform an additional set of fits using the error arrays to inject noise, but no emission line flux, to evaluate how often we recover a false positive significant line. We note that there is a false positive 1% of the time for 34605, while the others do not show false “detections” across 300 mocks with uncorrelated noise. However, 2% of the mocks for 23604 and 10155 yield C IV measurements, further supporting our use of a cutoff for C IV.
Appendix B Rest-Frame UV Fits to Line Emitters
Here, we present the emitters with their MultiNest fixed- and free-redshift fits. We also include some details about each of the AGN candidates.
B.0.1 10646 2
10646 (z=8.51) seems to be two objects, resolved in the UV (Weaver et al., in prep). One component, which we denote 10646 2, is a C IV, He II, and C III] emitter, while the other is included only as a C III] emitter, although it has C IV at the 3 level and evidence for He II+O III]. Neither one has compelling evidence for N IV] or N III]. There is a hint of [Ne V] in the combined spectrum, but this feature is also consistent with noise. All in all, 10646 2 is one of the strongest AGN candidates and also the highest-redshift among both the He II and C III] emitters in the sample. Fujimoto et al. (2023b) mentioned this object as a potential AGN with a nearby broad-line AGN (MSA ID 20466), both of which are seemingly surrounded by an ionization bubble, as evidenced by emission blueward of Ly.
B.0.2 45092
45092 (z=3.46, Bezanson et al. in prep) is a strong AGN candidate with C IV, He II+O III], Si III]+C III], and N IV]. Crucially, 45092 has the clearest, strongest [Ne V] of any source in the parent sample and is detected by Chandra (Bezanson et al. in prep). It is also one of the most massive galaxies in UNCOVER with .
B.0.3 47584
47584 appears to have He II(+O III]), N III], and (Si III]+)C III] but no C IV. Because of the N III], the lack of C IV, and the fact that 47584 sits at high sSFR (Figure 10), we must consider that this object is actually an exceptional low-metallicity nitrogen-enhanced SFG.
B.0.4 34605
We include 34605 because of its He II detection, but it is a questionable candidate because of the marginal detection (Section A), the lack of strong evidence for any other lines, and the overall noisiness of the spectrum.
B.0.5 BLAGN
LRD 45924 has He II emission (along with a host of other lines) but its line ratios place it in the star-forming region of the UV ratio diagnostics, as is expected for broad lines (Figure 6). As mentioned in Greene et al. (2024), 45924 is the one LRD with decent evidence of [Ne V].
We also find 11254 (z=6.87) has a broad H component, making it another strong AGN candidate. The residuals from a narrow-only fit illustrate the need for this component (Figure 14). It also joins 45924 in being more compact than all the other line emitters.